OpenStudy (anonymous):

The maximum value of 1cosx + 2cos2x + 3cos3x + ..... + 99cos99x

OpenStudy (anonymous):

The problem is that each cos is different

ganeshie8 (ganeshie8):

Hint : cos(0)

OpenStudy (anonymous):

yeah max is 1

OpenStudy (ikram002p):

any Fourier series :D if it goes to infinity it would be much easy xD it give the exact but here is a solution :- 0<=cos theta <=1

OpenStudy (anonymous):

but cos x = 1 for cos x for cos 3x = cos 0 so 3x =0 x =0 , fuk i tried it , it forms an ap yeah yeah

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

i should have contiued further

OpenStudy (ikram002p):

lol , but if x=0 then all goes to 1 :P so when u have x=0 it give u the maximum value :D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

all goes 1 ?

OpenStudy (ikram002p):

so no problem :3

OpenStudy (ikram002p):

yeb cos 0 = cos 2(0)= cos n (0) =1 for n in N :)

OpenStudy (ikram002p):

i guess there was a theorem , which i cant remember or idk if its true sum_n=0 ^ infinity n cos nx =1 for pi>n>0 hmm but not sure ( dont consider this case ) im only thinking in a comment :)

OpenStudy (anonymous):

i see

OpenStudy (ikram002p):

so now u know 1cosx + 2cos2x + 3cos3x + ..... + 99cos99x <=1+2+3....+99 =n(n+1)/2 for n=99

OpenStudy (kainui):

This is what it looks like when Gauss and Fourier hang out. =P

ganeshie8 (ganeshie8):

nice one ;)

OpenStudy (ikram002p):

yeah lol exactly xD